TOPOLOGY AND BASICITY IN POLYOXOVANADATES: AB INITIO STUDY OF THE CAGED MOLECULES [V_{18}O_{42}]^{12−} AND [V_{7}O_{12}(O_{3}PH)_{6}]^{−}

\underline{J. -P. BLAUDEAU}, M. -M. ROHMER, M. BÉNARD,, Laboratoire de Chimie Quantique, UPR 139 du CNRS, Université Louis Pasteur, 67000 Strasbourg, France; N. -E. GHERMANI, Laboratoire de Cristallographie et Modélisation des Matéraux Minéraux et Biologiques, CCM^{3}B, URA 809 du CNRS, Université Henri Poincaré-Nancy, 54506 Vandoeuvre-lès-Nancy, France.

Ab initio Hartree-Fock calculations have been carried out on the octadecavanadate ion [V_{18}O_{42}]^{12−}(1), on its protonated derivative [H_{4}V_{18}O_{42}]^{8−}(2), and on the vanadophosphate cluster [V_{7}O_{12}(O_{3}PH)_{6}]^{−}(3), taken as a model for [V_{7}O_{12}(O_{3}PR)_{6}]^{−} (3′). The spheroidal clusters (2) and (3′) have been recently characterized as “electronically inverse hosts” in the encapsulation complexes Cs_{9}[X@H_{4}V_{18}O_{42}]·12H_{2}O (X = Br, I) and (Ph_{4}P)_{2}[Cl@V_{7}O_{12}(O_{3}PPh)_{6}]. An estimate of the electrostatic potential distribution including the contribution of the lattice potential yields highly positive values of the potential inside the host cavities for (2) and (3). Those values are larger than the electrostatic potential computed at a vacant chloride site of the CsCl crystal, thus explaining the thermodynamic stability of the encapsulated anions.